# size of isogeny classes of abelian varieties in the extended lubin-tate case

### v1

@@ 3. Injectivity of the special fibre map
Let $\operatorname{End}(\mathscr{A})=R_{1}$ and $\operatorname{End}(A)=R_{2}$. We know that taking the special fibre is a fully faithful functor from the category of abelian varieties over $W(k)$ to abelian varieties over $k.$ Therefore we have a containment of orders $R_{1}\subset R_{2}$ inside $K$. We need to show that they are equal. Since $\widetilde{A}$ is a lift of $A$ to $L^{\prime}$, by Grothendieck- Messing theory, this corresponds to a lift of the filtration $\ker F\operatorname{mod}p\subset{\mathbb{D}}\operatorname{mod}p$ where ${\mathbb{D}}$ is the Dieudonné module of $A[p^{\infty}]$ and $F$ is the Frobenius. We note that all the endomorphisms of $A$ that lift to $\widetilde{A}$ are the ones that leave $\ker F$ invariant. Let $\widetilde{\beta}\in R_{2}.$ Since $R_{1}$ and $R_{2}$ are orders in $K$, their quotient is finite. Thus $n\widetilde{\beta}\in R_{1}$ for some $n.$ This implies that $\ker F$ is invariant under $n\widetilde{\beta}$. Note that $\ker F$ is invariant under the action of multiplication-by-$n$ map as it acts by multiplying its elements by $n$. This proves that $\widetilde{\beta}\in R_{1}.$
3.0.1. Outline of the proof of Proposition 3.1
Since $\widetilde{A}$ and $\widetilde{B}$ have the same CM type, we know that they are isogenous over ${\mathbb{C}}.$ In fact, by the Lemma 3.2, we have an isogeny $\widetilde{\alpha}:\widetilde{A}\rightarrow\widetilde{B}$ over $L^{\prime}.$ The Néron mapping theorem extends this to a unique $W(k)$-morphism of schemes $\varphi:\mathscr{A}\rightarrow\mathscr{B}$. The map $\varphi$ induces an endomorphism $\alpha$ of the special fibre of $A$, that is $\alpha:A\rightarrow A.$ Moreover, by Lemma 3.3, this lifts to an endomorphism $\psi$ of $\mathscr{A},$ generically which gives us an endomorphism $\widetilde{\beta}:\widetilde{A}\rightarrow\widetilde{A}$. Let $G_{\varphi}$ and $G_{\psi}$ be the kernels of $\phi$ and $\psi$ respectively. These are finite flat group schemes over $W(k)$ that reduce to the kernel of $\alpha$ mod $v$. We embed them into an $N$-torsion group scheme $\mathscr{A}[n]$ which is the kernel of the multiplication by $n$ map on $\mathscr{A}$ for some positive integer $n.$ We will prove that the kernels of the two maps must agree as group schemes. Moreover, since we have the same group scheme that occurs as the kernel of an isogeny between $\widetilde{A}$ and $\widetilde{B}$ and an endomorphism of $\widetilde{A}$, we deduce that $\widetilde{A}$ is in fact isomorphic to $\widetilde{B}.$ To that end we have the following Lemma.
Lemma 3.4.
Let $G_{\varphi}$, $G_{\psi}$ and $\mathscr{A}[n]$ be as described above. Let $H$ denote the special fibre of $G_{\varphi}$, $G_{\psi}\subset A[n]$. Then the group subgroup-scheme $G_{\varphi}$ equals $G_{\psi}$.
In order to prove Lemma 3.4, we will use Fontaine’s results on the classification of finite flat group schemes of $p$-power order over an unramified base. For reference, we state the results that we use below from [Con99]. We note that these result only holds for odd primes and we therefore restrict to that case as well for the results in this paper.
3.1. Classification results for finite flat group schemes of prime power order over an unramified base.
@@
The category of finite Honda systems over $k$ consists of tuples $(\mathscr{L},\mathscr{M})$ where $\mathscr{L}$ and $\mathscr{M}$ are $W(k)[F,V]$-modules of finite $W(k)-$ length such that the following is satisfied.
@@
    The map $\mathscr{L}/p\mathscr{L}\rightarrow F\mathscr{M}/p\mathscr{M}$ is an isomorphism.
@@
(Fontaine, see [Con99] Lemma 1.3) The category of finite Honda systems over $k$ is abelian. For a morphism $\varphi:(\mathscr{L},\mathscr{M})\rightarrow(\mathscr{L},\mathscr{M})$, the kernel $\ker(\varphi)=(\mathscr{L}^{\prime},\mathscr{M}^{\prime})$ where $\mathscr{M}^{\prime}=\ker(\mathscr{M}\rightarrow\mathscr{M})$ and $\mathscr{L}^{\prime}=\mathscr{L}\cap\mathscr{M}^{\prime}$ and $\operatorname{coker}(\varphi)=(\mathscr{L}^{\prime\prime},\mathscr{M}^{\prime\prime})$ where ′′$\mathscr{M}^{\prime\prime}=\operatorname{coker}(\mathscr{M}\rightarrow\mathscr{M})$ and $\mathscr{L}^{\prime\prime}$ is the image of the composition of maps of $W(k)$-modules $\mathscr{L}\rightarrow\mathscr{M}\rightarrow\mathscr{M}^{\prime\prime}.$
Theorem 3.7.
@@
(Fontaine, see Theorem 1.4 [Con99]) For $p>2$, there is an anti-equivalence between the category of finite flat commutative group schemes $G$ over $W(k)$ of $p$-power order and the category of finite Honda systems over $k$. Under this equivalence, $G$ maps to $(\mathscr{L},\mathscr{M})$ where $\mathscr{M}$ is Dieudonné module of the special fibre $G_{k}$ and $\mathscr{L}$ is a $W(k)$ sub-module of $\mathscr{M}$ satisfying the three properties in Definition 3.5. See, for example [Con99], page $5-6$ for a complete description.
The proof of Lemma 3.4 will now follow by proving the result in the category of finite Honda systems as follows.
Proof of Lemma 3.4.
@@
Since all the group schemes are embedded inside the kernel of the multiplication map $[n]$, it is enough to consider the cases of group schemes with prime to $p$ order and $p$-power order separately. For $(n,p)=1$, the reduction map of an abelian variety over $L^{\prime}$ to its special fibre induces an isomorphism on the $n$-torsion. Therefore, the prime-to-$p$ subgroups lift uniquely. Moving on to the case where $n=p^{r}$, under the anti-equivalence in Proposition 3.7, we associate the tuples $(\mathscr{L},\mathscr{M}),(\mathscr{L}_{1},\mathscr{M}_{1})$ and $(\mathscr{L}_{2},\mathscr{M}_{2})$ to the group schemes $\mathscr{A}[p^{r}],G_{\varphi}$ and $G_{\psi}$ respectively. Since both the kernels reduce to the same group scheme over $k,$ we see that $\mathscr{M}_{1}=\mathscr{M}_{2}$ as they correspond to the Dieudonné module of the special fibre $T$. Furthermore, the tuples $(\mathscr{L}_{1},\mathscr{M}_{1})$ and $(\mathscr{L}_{2},\mathscr{M}_{2})$ are co-kernels of $(\mathscr{L},\mathscr{M})$ in the category of finite Honda systems. Therefore, Proposition 3.6 proves that $\mathscr{L}_{1}=\mathscr{L}_{2}.$ This shows that the group schemes $G_{\varphi}$ and $G_{\psi}$ are the same as subgroup-schemes of $\mathscr{A}[p^{r}]$ and hence, the required kernels are the same. ∎
Finally, the proposition follows from the following argument that we give below.
Proof of Proposition 3.1.
@@
Lemma 3.4 shows that $\ker\widetilde{\alpha}=\ker\widetilde{\beta}$. This gives us maps $\widetilde{\varphi}:\widetilde{A}\rightarrow\widetilde{B}$ and $\widetilde{\psi}:\widetilde{A}\rightarrow\widetilde{B}$ such that $\widetilde{\varphi}\circ\widetilde{\alpha}=\widetilde{\beta}$ and $\widetilde{\psi}\circ\widetilde{\beta}=\widetilde{\alpha}$. Then we can check that $\widetilde{\varphi}\circ\widetilde{\psi}$ and $\widetilde{\psi}\circ\widetilde{\varphi}$ act as identity when we compose them with $\widetilde{\beta}$ and $\widetilde{\alpha}$ respectively, thus completing the proof of proposition.
4. Local conditions for the smallest endomorphism ring
@@
We next find the local conditions of the orders that can occur as an endomorphism ring in the isogeny class of $A$ over ${\mathbb{F}}_{q}.$ We recall that the Newton polygon of $A[p^{\infty}]$ has slope $1/g$ and $(g-1)/g$. Moreover, we assume that the prime $p$ splits in ${\mathcal{O}}_{K}$ as ${\mathfrak{p}}_{1}$ and ${\mathfrak{p}}_{2}.$ Therefore, the endomorphism algebra $K$ of $A$ satisfies the condition that $K\otimes{\mathbb{Q}}_{p}=K_{1}\oplus K_{2}$ where $K_{1}$ and $K_{2}$ are unramified extensions of ${\mathbb{Q}}_{p}$ of degree $g$.
@@ 4.1. Local conditions at the prime $p$
Let $A$ be an abelian variety over ${\mathbb{F}}_{q}$ with endomorphism algebra $K$ as described above. Let $\operatorname{End}(A)=R.$ Then the order $R$ is maximal at $p$. More precisely, $R\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}}$ where ${\mathcal{O}}_{K_{1}}$ (respectively ${\mathcal{O}}_{K_{2}}$) are the maximal orders in $K_{1}$ (respectively $K_{2}$).
@@
Let ${\mathbb{D}}$ denote the Dieudonné module of $A[p^{\infty}]$. Note that ${\mathbb{D}}$ admits a decomposition locally as a direct sum ${\mathbb{D}}\otimes{\mathbb{Q}}_{p}={\mathbb{D}}_{1}\times{\mathbb{D}}_{2}$ where ${\mathbb{D}}_{1}$ and ${\mathbb{D}}_{2}$ are modules over $K_{{\mathfrak{p}}_{1}}$ and $K_{{\mathfrak{p}}_{2}}$ respectively. This further gives us a decomposition of the $p$-divisible group $A[p^{\infty}]\simeq{\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$ over ${\mathbb{F}}_{q}$ where ${\mathscr{G}}_{1}$ and ${\mathscr{G}}_{2}$ have slopes $1/g$ and $(g-1)/g$ respectively (see Theorem 9.2 in [Oor07]). Since ${\mathscr{G}}_{1}$ cannot be isogenous to ${\mathscr{G}}_{2}$, we see that $\operatorname{End}({\mathscr{G}}_{1}\times{\mathscr{G}}_{2})\simeq\operatorname{End}({\mathscr{G}}_{1})\times\operatorname{End}({\mathscr{G}}_{2})$. We know that Tate’s theorem states that $R\otimes{\mathbb{Z}}_{p}\simeq\operatorname{End}A[p^{\infty}]$. Therefore, it is sufficient to prove that $\operatorname{End}({\mathscr{G}}_{1})$ and $\operatorname{End}({\mathscr{G}}_{2})$ is maximal. This can be shown analogously to an argument in [OS20], Proposition $2.1$ which we give below. Let ${\mathscr{G}}^{\prime}$ be a $p$-divisible group isogenous to ${\mathscr{G}}_{1}$ such that $\operatorname{End}({\mathscr{G}}^{\prime})$ is maximal in its field of fractions. Let $j:{\mathscr{G}}_{0}\rightarrow{\mathscr{G}}_{1}$ be any arbitrary isogeny. Since ${\mathscr{G}}_{1}$ is isomorphic to a quotient of ${\mathscr{G}}_{0}$ by the kernel of $j$, we will show that $\operatorname{End}({\mathscr{G}}^{\prime})$ preserves $\ker j,$ proving that $\operatorname{End}({\mathscr{G}}^{\prime})\subset\operatorname{End}({\mathscr{G}}_{1}).$ Since ${\mathscr{G}}_{0}$ is a one dimensional $p$-divisible group, it has a unique subgroup of order $p^{r}$ for each $r\geq 1$. Therefore, $\operatorname{End}({\mathscr{G}}^{\prime})$ preserves the kernel of $j$. The same argument with the dual of ${\mathscr{G}}_{1}$ proves that $\operatorname{End}({\mathscr{G}}_{2})$ is maximal. ∎

### v2

@@ 2. Injectivity of the special fibre map
Let $\widetilde{\mathscr{A}}$ be the Néron model of $\widetilde{A}$. We denote by $\operatorname{End}(\widetilde{A})=\operatorname{End}(\widetilde{\mathscr{A}})=R$ and $\operatorname{End}(A)=S$. We know that taking the special fibre is a fully faithful functor from the category of abelian varieties over $W(k)$ to abelian varieties over $k.$ Therefore we have a containment of orders $R\subset S$ inside $K$. We need to show that they are equal. Since $\widetilde{A}$ is a lift of $A$ to $L^{\prime}$, by Grothendieck- Messing theory, this corresponds to a lift of the filtration $\ker F\operatorname{mod}p\subset{\mathbb{D}}\operatorname{mod}p$ where ${\mathbb{D}}$ is the Dieudonné module of $A[p^{\infty}]$ and $F$ is the Frobenius. We note that all the endomorphisms of $A$ that lift to $\widetilde{A}$ are the ones that leave $\ker F$ invariant. Let $\widetilde{\beta}\in S.$ Since $R$ and $S$ are orders in $K$, their quotient is finite. Thus $n\widetilde{\beta}\in R$ for some $n.$ This implies that $\ker F$ is invariant under $n\widetilde{\beta}$. Note that $\ker F$ is invariant under the action of multiplication-by-$n$ map as it acts by multiplying its elements by $n$. This proves that $\widetilde{\beta}\in R.$
The main result of this section is Proposition 2.7 which provides an injection between the isogeny class of $\widetilde{A}$ over $L^{\prime}$ to the isogeny class of its reduction mod $p$. The reader may wish to skip to the proposition directly. Its proof involves Fontaine’s results on classification of finite flat group schemes that we first review below.
2.1. Classification of finite flat group schemes of prime power order.
We follow the exposition and results in [Con99] Section 1, and state the results that we apply later. We work over an unramified base and denote by $D_{k}:=W(k)[F,V]$ where $F$ and $V$ will stand for the Frobenius and the Verchiebung. We recall that $p>2$.
@@
The category of finite Honda systems over $W(k)$ consists of tuples $(\mathscr{L},\mathscr{M})$ where $\mathscr{L}$ and $\mathscr{M}$ are $D_{k}$-modules of finite $W(k)$-length such that the following is satisfied:
@@
    The $k$-linear map $\mathscr{L}/p\mathscr{L}\rightarrow\mathscr{M}/F\mathscr{M}$ is an isomorphism.
@@
(Fontaine. See [Con99] Lemma 1.3) The category of finite Honda systems over $k$ is abelian. Let $\varphi:(\mathscr{L},\mathscr{M})\rightarrow(\mathscr{L},\mathscr{M})$ be a morphism in this category. Then the kernel is given by $\ker(\varphi):=(\mathscr{L}^{\prime},\mathscr{M}^{\prime})$ where $\mathscr{M}^{\prime}:=\ker(\mathscr{M}\rightarrow\mathscr{M})$ and $\mathscr{L}^{\prime}:=\mathscr{L}\cap\mathscr{M}^{\prime}$. While the cokernel is $\operatorname{coker}(\varphi):=(\mathscr{L}^{\prime\prime},\mathscr{M}^{\prime\prime})$ where ′′$\mathscr{M}^{\prime\prime}:=\operatorname{coker}(\mathscr{M}\rightarrow\mathscr{M})$ and $\mathscr{L}^{\prime\prime}$ is the image of the composition of maps of $W(k)$-modules $\mathscr{L}\rightarrow\mathscr{M}\rightarrow\mathscr{M}^{\prime\prime}.$
Proposition 2.5.
@@
(Fontaine. See Theorem 1.4 [Con99]) For $p>2$, there is an anti-equivalence between the category of finite flat commutative group schemes $G$ over $W(k)$ of $p$-power order and the category of finite Honda systems over $W(k)$.
Under this equivalence, $G$ maps to $(\mathscr{L},\mathscr{M})$ where $\mathscr{M}$ is Dieudonné module of the special fiber which we denote by $G_{k}$ and $\mathscr{L}$ is a $W(k)$ sub-module of $\mathscr{M}$ satisfying the three properties in Definition 2.3. See, for example [Con99], Section $1$ for a summary of Fontaine’s results.
Now, let $\widetilde{A}$ and $\widetilde{B}$ be as in Lemma 2.1. Since $\widetilde{A}$ and $\widetilde{B}$ have the same CM type, by Lemma 2.1, we have an isogeny $\widetilde{\alpha}:\widetilde{A}\rightarrow\widetilde{B}$ over $L^{\prime}.$ The Néron mapping theorem extends this to a unique $W(k)$-morphism between their Néron models $\varphi:\widetilde{\mathscr{A}}\rightarrow\widetilde{\mathscr{B}}$. The map $\varphi$ induces an endomorphism $\alpha$ of the special fibre of $A$, that is $\alpha:A\rightarrow A.$ Moreover, by Lemma 2.2, this lifts to an endomorphism $\psi$ of $\mathscr{A},$ generically which gives us an endomorphism $\widetilde{\beta}:\widetilde{A}\rightarrow\widetilde{A}$. We will need the following lemma.
Lemma 2.6.
We use the notation above. Let $G_{\varphi}$ and $G_{\psi}$ be the kernels of $\varphi$ and $\psi$ respectively. For some $m>0$, we have $G_{\varphi},G_{\psi}\subset\widetilde{\mathscr{A}}[m]$. Let $H\subset A[m]$ denote the special fibre of the two finite flat group schemes. Then $G_{\varphi}$ equals $G_{\psi}$ as subgroup schemes of $\widetilde{\mathscr{A}}[m]$.
Proof.
@@
Let $G_{\varphi}$, $G_{\psi}$ and $\widetilde{\mathscr{A}}[m]$ be as described above. Let $H$ denote the special fibre of $G_{\varphi}$, $G_{\psi}\subset A[m]$. Since all the group schemes are embedded inside the kernel of the multiplication map $[m]$, it is enough to consider the cases of group schemes with prime to $p$ order and $p$-power order separately. For $(m,p)=1$, the reduction map of an abelian variety over $L^{\prime}$ to its special fibre induces an isomorphism on the $m$-torsion. Therefore, the prime-to-$p$ subgroups lift uniquely. For the case where $m=p^{r}$, we associate the tuples $(\mathscr{L},\mathscr{M}),(\mathscr{L}_{1},\mathscr{M}_{1})$ and $(\mathscr{L}_{2},\mathscr{M}_{2})$ to the group schemes $\widetilde{\mathscr{A}}[p^{r}],G_{\varphi}$ and $G_{\psi}$ respectively, under the anti-equivalence in Proposition 2.5. Since both the kernels reduce to the same group scheme over $k,$ we see that $\mathscr{M}_{1}=\mathscr{M}_{2}$ as they correspond to the Dieudonné module of the special fibre $T$. Furthermore, the tuples $(\mathscr{L}_{1},\mathscr{M}_{1})$ and $(\mathscr{L}_{2},\mathscr{M}_{2})$ are co-kernels of $(\mathscr{L},\mathscr{M})$ in the category of finite Honda systems. Therefore, Proposition 2.4 proves that $\mathscr{L}_{1}=\mathscr{L}_{2}.$ This shows that the group schemes $G_{\varphi}$ and $G_{\psi}$ are the same as subgroup-schemes of $\widetilde{\mathscr{A}}[p^{r}]$ and hence, the required kernels are the same. ∎
We end this section by proving the following proposition.
Proposition 2.7.
Let $\widetilde{A}$ and $\widetilde{B}$ be as in Lemma 2.1. Let $\widetilde{\mathscr{A}}$ and $\widetilde{\mathscr{B}}$ denote their Néron models respectively over $W(k)$ such that their special fibre is is $A.$ Then $\widetilde{A}$ and $\widetilde{B}$ are isomorphic.
Proof.
@@
We recall the notation of $\widetilde{\alpha}$ and $\widetilde{\beta}$ mentioned before Lemma 2.6. We know that $\ker\widetilde{\alpha}=\ker\widetilde{\beta}$ by Lemma 2.6 . Therefore, we get maps $\widetilde{\varphi}:\widetilde{A}\rightarrow\widetilde{B}$ and $\widetilde{\psi}:\widetilde{A}\rightarrow\widetilde{B}$ such that $\widetilde{\varphi}\circ\widetilde{\alpha}=\widetilde{\beta}$ and $\widetilde{\psi}\circ\widetilde{\beta}=\widetilde{\alpha}$. Then we can check that $\widetilde{\varphi}\circ\widetilde{\psi}$ and $\widetilde{\psi}\circ\widetilde{\varphi}$ act as identity when we compose them with $\widetilde{\beta}$ and $\widetilde{\alpha}$ respectively, thus completing the proof of proposition. ∎
3. Local conditions for the endomorphism ring
@@
We recall that the Newton polygon of $A[p^{\infty}]$ has slope $1/g$ and $(g-1)/g$. Moreover, we assume that the prime $p$ splits in ${\mathcal{O}}_{K}$ such that $K\otimes{\mathbb{Q}}_{p}=K^{\prime}\oplus K^{\prime}$ where $K^{\prime}$ is the unramified extensions of ${\mathbb{Q}}_{p}$ of degree $g$. We next find the local conditions of the orders that can occur as an endomorphism ring in the isogeny class of $A$ over ${\mathbb{F}}_{q}.$
@@ 3.1. Local conditions at the prime $p$
Let $A$ be an abelian variety over ${\mathbb{F}}_{q}$ with endomorphism algebra $K={\mathbb{Q}}(\alpha)$ where $\alpha$ is the Weil-$q$ number of $A$. Let $\operatorname{End}(A)=R.$ Then the order $R$ is maximal at $p$. More precisely, $R\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$ where ${\mathcal{O}}_{K^{\prime}}$ is the maximal order in $K^{\prime}$.
@@
Let ${\mathbb{D}}$ denote the Dieudonné module of $A[p^{\infty}]$. Note that ${\mathbb{D}}$ admits a decomposition locally as a direct sum ${\mathbb{D}}\otimes{\mathbb{Q}}_{p}={\mathbb{D}}_{1}\times{\mathbb{D}}_{2}$ where ${\mathbb{D}}_{1}$ and ${\mathbb{D}}_{2}$ are modules over $K^{\prime}$. This further gives us a decomposition of the $p$-divisible group $A[p^{\infty}]\simeq{\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$ over ${\mathbb{F}}_{q}$ where ${\mathscr{G}}_{1}$ and ${\mathscr{G}}_{2}$ have slopes $1/g$ and $(g-1)/g$ respectively (see Theorem 9.2 in [Oor08]), and ${\mathscr{G}}_{2}$ is the dual of ${\mathscr{G}}_{1}$. Since ${\mathscr{G}}_{1}$ cannot be isogenous to ${\mathscr{G}}_{2}$, we see that $\operatorname{End}({\mathscr{G}}_{1}\times{\mathscr{G}}_{2})\simeq\operatorname{End}({\mathscr{G}}_{1})\times\operatorname{End}({\mathscr{G}}_{2})$. We know that Tate’s theorem states that $R\otimes{\mathbb{Z}}_{p}\simeq\operatorname{End}A[p^{\infty}]$. Therefore, it is sufficient to prove that $\operatorname{End}({\mathscr{G}}_{1})$ and $\operatorname{End}({\mathscr{G}}_{2})$ are maximal. This argued as in [OS20], Proposition $2.1$ which we describe below. Let ${\mathscr{G}}^{\prime}$ be a $p$-divisible group isogenous to ${\mathscr{G}}_{1}$ such that $\operatorname{End}({\mathscr{G}}^{\prime})$ is maximal in its field of fractions. Let $j:{\mathscr{G}}_{0}\rightarrow{\mathscr{G}}_{1}$ be any arbitrary isogeny. Since ${\mathscr{G}}_{1}$ is isomorphic to a quotient of ${\mathscr{G}}_{0}$ by the kernel of $j$, we will show that $\operatorname{End}({\mathscr{G}}^{\prime})$ preserves $\ker j,$ proving that $\operatorname{End}({\mathscr{G}}^{\prime})\subseteq\operatorname{End}({\mathscr{G}}_{1}),$ proving that $\operatorname{End}({\mathscr{G}}_{1})$ is maximal. Since ${\mathscr{G}}_{0}$ is a one dimensional $p$-divisible group, it has a unique subgroup of order $p^{r}$ for each $r\geq 1$. Therefore, $\operatorname{End}({\mathscr{G}}^{\prime})$ preserves the kernel of $j$. The same argument with the dual of ${\mathscr{G}}_{1}$ proves that $\operatorname{End}({\mathscr{G}}_{2})$ is maximal. ∎